# Deep Learning Meets Mechanism Design: A Survey

### v2

@@ VI-2 Formulation 2
Figure 15: NN for Groves mechanism payment computation for each buyer $i$ (Figure reproduced from [AHMAD21])
@@ VI-3 Formulation 3
[MISHRA22] introduces the concept of EEF1 Allocation (efficient and envy-free up to one item). The formulation is as follows. Consider a set $M=[m]$ of indivisible items to be allocated among $N=[n]$ buyers. Each agent $i\in N$ has a valuation function $v_{i}:2^{M}\rightarrow R$. The valuation profile is represented as $(v_{1},v_{2},...,v_{n})$ and the valuations are additive. The valuation of a buyer $i$ for a bundle $A_{i}$ is $v_{i}(A_{i})=\sum_{j\in A_{i}}v_{i}(\{j\})$. The Utilitarian Social Welfare (USW) is defined as $sw(A,v)=\sum_{i\in N}v_{i}(A_{i})$ and $v_{i}$ is assumed to be drawn from known distributions for each agent. The output allocation is $A\in\{0,1\}^{n\text{x}m}$ where $A_{i}\in[m]$ is the bundle assigned to buyer $i$. $A_{i}\bigcap A_{k}=\phi\;\forall i,k\in N$ and $\bigcup_{i}A_{i}=M$.
@@
Figure 16: EEF1-NN Architecture (Figure reproduced from [MISHRA22])
The NN architecture used in [MISHRA22] is shown in Fig. 16. The optimization problem is formulated as
@@ VII Learning Mechanisms with Budget Balance
Auctions are required in some situations to purely elicit true valuations from the participants, for example, to allocate public goods where no money needs to be collected. [MANISHA18] designs a neural network, for such situations, to redistribute the payments made back to the participants such that DSIC is not compromised. It considers a setting where $p$ public goods need to be allocated to $n$ competing agents where each agent wants at most one of these goods. If the goods are homogeneous, agent $i$ values each good at $v_{i}=\theta_{i}$ whereas if they are heterogeneous, the valuations for each good for agent $i$ is given as $v_{i}=(\theta_{i1},\theta_{i2},...,\theta_{ip})$. Allocative efficiency is ensured by calculating the allocations analytically as a function of the submitted bids.
@@
Figure 17: NN architecture for linear rebate function (Figure reproduced from [MANISHA18])
Figure 18: NN architecture for nonlinear rebate function (Figure reproduced from [MANISHA18])
@@ VIII Three Illustrative Applications
Deep learning as a technique to design complex mechanisms has been explored in many applications. Luong et. al. [LUONG18] design an optimal auction in the context of edge computing resource management in mobile networks. Luong and co-authors [LUONG23] look into an interesting energy management application in vehicular networks using MyersonNet. Qian et. al. [QIAN19] have applied the approach to resource allocation in wireless virtualisation. Bharadwaj and co-authors [BHARDWAJ23] have explored an interesting input procurement application in agriculture for designing a versatile auction that satisfies incentive compatibility, individual rationality, cost minimization fairness of allocation, and certain business constraints. Zhou et. al. [ZHOU18] have used the approach for predicting clickthrough rate in internet advertising while Liu et. al. [LIU21] have explored the approach in e-commerce advertising. Golowich and co-authors [GOLOWICH18] have used this approach for the multifacility location problem which arises in numerous contexts. Cepeda [Cepeda] has investigated this approach in the context of electricity auctions where the author has extended the technique in Dutting et al. [DUTTING23] to discover nearly optimal designs for electricity auctions incorporating capacity constraints and realistic assumptions.
We now discuss in some detail, the applications discussed by Luong et. al. [LUONG18], Qian et. al. [QIAN19], and Bhardwaj et. al. [BHARDWAJ23].
@@ VIII-A Efficient Energy Management in a UAV Assisted Vehicular Network
In [LUONG23], the authors study efficient energy management using a neural network enabled auction mechanism. In this example, multiple unmanned aerial vehicles (UAVs) are deployed to collect data related to physical entities, such as vehicle information and roadway infrastructure. The collected data are then communicated; this ensures data synchronization with a virtual service provider (VSP). The VSP creates, using this data, digital twins of the physical entities. Metaverse platform is used to create digital twins. Fig. 19. As shown in the figure, each UAV returns to its base after gathering data, transmits this information to the VSP and charges itself via an Energy Service Provider (ESP). The ESP may contain one or more Mobile Charging Stations (MCSs). As the capacity of MCSs is limited and there is a revenue requirement for the ESP, the allocation of the UAVs must be done keeping these constraints in mind. One way of solving this energy management problem is to use traditional optimization methods with a set of constraints. However, this does not consider economic factors, such as profit and loss. The authors use a mechanism design approach by exploring auction based mechanisms that consider the profit, cost, and revenue.
@@
Figure 19: (a) A UAV assisted vehicular metaverse synchronisation system (b) A deep learning based auction model (Figure reproduced from[LUONG23])
@@
Figure 20: Feed forward neural network for allocation and payment rules (Figure reproduced from [LUONG23])
It is observed in [LUONG23], through simulations, that the proposed auction based scheme outperforms the classical second price auction in terms of revenue.
@@ VIII-B Resource Allocation in a Mobile Network
A Mobile Network Virtual Operator (MNVO) provides services to its customers who do not own any wireless infrastructure. A wireless virtualization enables MNVO to use multiple virtual wireless networks to provide services over the same physical infrastructure. The allocation of resources owned by the MNVO to its mobile customers is a challenging problem. The objective of MNVOs is to allocate resources (sub-channel and power) to maximize their respective revenues. Classical auction methods typically maximize social welfare and may often perform poorly in terms of revenues. In [QIAN19], the authors present a deep learning network to learn an auction-based method to allocate resources to the users of the MNVo. Fig. 21 shows the proposed auction model with the MNVO auctioning the resources (sub-channel and power) to a set of $N=\{1,\dots,n\}$ users. Let user $i$ get power $p_{i}$ and sub-channel $sc_{i}$. If these values are higher than the corresponding values for other users $j\neq i$, the valuation of the $i^{th}$ user is given as $v_{i}=p_{i}\times sc_{i}$. Let the payment for user $i$ be $t_{i}$, then the utility is
Figure 21: Auction for allocating sub-channels and power (Figure reproduced from [QIAN19])
@@
The optimization problem of Eqs. 98-100 is solved using MyersonNet [DUTTING23] given in Figs. 5,6,7. It is observed that the deep learning based resource allocation improves the revenue for MNVO as compared to Second Price Auction (with traditional optimization).
@@ VIII-C A Procurement Auction for Agricultural Inputs
[BHARDWAJ23] extends the RegretNet architecture [DUTTING23] to tackle the problem of volume discount procurement auctions in the agricultural setting. Farmer Producer Organisations (FPOs) or Farmer Collectives (FCs) typically consist of around thousand farmers. Individual farmers may have requirements such as seeds, pesticides, fertilizers, etc., but may not be able to bargain the best deal from the suppliers if they approach the suppliers on their own. FPOs collect the requirements of all their farmers, organize all the requirements into clusters, and approach the suppliers with a bulk requirement. Due to the large volume of procurement, suppliers generally offer volume discounts, i.e., a higher percentage of discount on the original asking price as the number of units purchased increases. This is reflected in the bids submitted by the suppliers. In addition to the constraints satisfied by the RegretNet architecture, namely, IR, DSIC, and revenue maximization, the authors aim to satisfy certain additional constraints. Revenue maximization here becomes cost minimization, as it is a procurement auction. The additional constraints are fairness (envy-freeness), social welfare maximization, and business constraints (minimum and maximum number of winning suppliers, minimum amount of business to be given to each winning supplier, etc.). There is, however, some compromise to the social welfare maximization.
@@
The volume discount procurement auction setting of [BHARDWAJ23] is as follows. Consider the case of a single buyer (FPO) who wishes to purchase $m$ homogeneous units of a single item from $n$ different suppliers. Let $m$ be divided into $k$ brackets. For each bracket, the bid submitted by the supplier (which is the price per unit in that bracket) must be less than or equal to the bid submitted by the supplier for the previous bracket. Each supplier $i$ submits a bid vector $b^{(i)}=(b_{1}^{(i)},b_{2}^{(i)},...,b_{k}^{(i)})$. Given all the input bids $b=(b^{(1)},b^{(2)},...,b^{(n)})$, the neural network outputs the allocation vector $a(b)=(a_{1}(b),a_{2}(b),...,a_{n}(b))$ and the payment vector $p(b)=(p_{1}(b),p_{2}(b),...,p_{n}(b))$ where $a_{i}(b)$ is the number of units sold by supplier $i$ and $p_{i}(b)$ is the payment made to supplier $i$.
@@
Through experimental analysis, the authors of [BHARDWAJ23] show how a DL based procurement auction captures the essence of the agricultural market better than existing analytical methods. Even as it minimizes additional constraints such as envy, when compared to VCG, their model saves the FPO more money.
@@ IX Summary and Conclusions
Table II provides a summary of the more prominent DL architectures for auction design as well as their evaluation. The DL approaches, in current literature, used for designing auctions include MenuNet, Deep Interest Network, RochetNet, MoulineNet, RegretNet, PreferenceNet, Deep Neural Auction and the Reinforcement Learning setups. MenuNet [SHEN18] is based on the principle of taxation, which lets buyers make selections that yield an IC mechanism. In online auctions, the essential task of click through rate prediction is implemented using Deep Interest Network [ZHOU18]. MoulineNet is used to determine agents’ preferred optimal facility location [GOLOWICH18]. An extension of RegretNet is PreferenceNet [PERI21] that includes buyers’ preferences in Auction design. The Reinforcement learning setup is based on a no-regret learning mechanism and in the Deep Neural Auction model [LIU21], the positional effect of ordering of sellers based on some metric is removed. The summary of key approaches referenced in our paper are given in Table II. In this table, strategyproofness indicates the type of IC constraint the model aims to satisfy, efficiency indicates whether it is revenue maximizing or welfare maximizing, fairness indicates the model’s aim to minimize violations of IR, envy, etc. and other constraints refer to constraints such as budget balance that the model aims to satisfy.

### v3

@@ VI-2 Formulation 2
Figure 15: NN for Groves mechanism payment computation for each buyer $i$ (Figure reproduced from [1])
@@ VI-3 Formulation 3
[32] introduces the concept of EEF1 Allocation (efficient and envy-free up to one item). The formulation is as follows. Consider a set $M=[m]$ of indivisible items to be allocated among $N=[n]$ buyers. Each agent $i\in N$ has a valuation function $v_{i}:2^{M}\rightarrow R$. The valuation profile is represented as $(v_{1},v_{2},...,v_{n})$ and the valuations are additive. The valuation of a buyer $i$ for a bundle $A_{i}$ is $v_{i}(A_{i})=\sum_{j\in A_{i}}v_{i}(\{j\})$. The Utilitarian Social Welfare (USW) is defined as $sw(A,v)=\sum_{i\in N}v_{i}(A_{i})$ and $v_{i}$ is assumed to be drawn from known distributions for each agent. The output allocation is $A\in\{0,1\}^{n\text{x}m}$ where $A_{i}\in[m]$ is the bundle assigned to buyer $i$. $A_{i}\bigcap A_{k}=\phi\;\forall i,k\in N$ and $\bigcup_{i}A_{i}=M$.
@@
Figure 16: EEF1-NN Architecture (Figure reproduced from [32])
The NN architecture used in [32] is shown in Fig. 16. The optimization problem is formulated as
@@ VII Learning Mechanisms with Budget Balance
Auctions are required in some situations to purely elicit true valuations from the participants, for example, to allocate public goods where no money needs to be collected. [29] designs a neural network, for such situations, to redistribute the payments made back to the participants such that DSIC is not compromised. It considers a setting where $p$ public goods need to be allocated to $n$ competing agents where each agent wants at most one of these goods. If the goods are homogeneous, agent $i$ values each good at $v_{i}=\theta_{i}$ whereas if they are heterogeneous, the valuations for each good for agent $i$ is given as $v_{i}=(\theta_{i1},\theta_{i2},...,\theta_{ip})$. Allocative efficiency is ensured by calculating the allocations analytically as a function of the submitted bids.
@@
Figure 17: NN architecture for linear rebate function (Figure reproduced from [29])
Figure 18: NN architecture for nonlinear rebate function (Figure reproduced from [29])
@@ VIII Three Illustrative Applications
Deep learning as a technique to design complex mechanisms has been explored in many applications. Luong et. al. [28] design an optimal auction in the context of edge computing resource management in mobile networks. Luong and co-authors [27] look into an interesting energy management application in vehicular networks using MyersonNet. Qian et. al. [39] have applied the approach to resource allocation in wireless virtualisation. Bharadwaj and co-authors [4] have explored an interesting input procurement application in agriculture for designing a versatile auction that satisfies incentive compatibility, individual rationality, cost minimization fairness of allocation, and certain business constraints. Zhou et. al. [49] have used the approach for predicting clickthrough rate in internet advertising while Liu et. al. [26] have explored the approach in e-commerce advertising. Golowich and co-authors [17] have used this approach for the multifacility location problem which arises in numerous contexts. Cepeda [8] has investigated this approach in the context of electricity auctions where the author has extended the technique in Dutting et al. [14] to discover nearly optimal designs for electricity auctions incorporating capacity constraints and realistic assumptions.
We now discuss in some detail, the applications discussed by Luong et. al. [28], Qian et. al. [39], and Bhardwaj et. al. [4].
@@ VIII-A Efficient Energy Management in a UAV Assisted Vehicular Network
In [27], the authors study efficient energy management using a neural network enabled auction mechanism. In this example, multiple unmanned aerial vehicles (UAVs) are deployed to collect data related to physical entities, such as vehicle information and roadway infrastructure. The collected data are then communicated; this ensures data synchronization with a virtual service provider (VSP). The VSP creates, using this data, digital twins of the physical entities. Metaverse platform is used to create digital twins. Fig. 19. As shown in the figure, each UAV returns to its base after gathering data, transmits this information to the VSP and charges itself via an Energy Service Provider (ESP). The ESP may contain one or more Mobile Charging Stations (MCSs). As the capacity of MCSs is limited and there is a revenue requirement for the ESP, the allocation of the UAVs must be done keeping these constraints in mind. One way of solving this energy management problem is to use traditional optimization methods with a set of constraints. However, this does not consider economic factors, such as profit and loss. The authors use a mechanism design approach by exploring auction based mechanisms that consider the profit, cost, and revenue.
@@
Figure 19: (a) A UAV assisted vehicular metaverse synchronisation system (b) A deep learning based auction model (Figure reproduced from[27])
@@
Figure 20: Feed forward neural network for allocation and payment rules (Figure reproduced from [27])
It is observed in [27], through simulations, that the proposed auction based scheme outperforms the classical second price auction in terms of revenue.
@@ VIII-B Resource Allocation in a Mobile Network
A Mobile Network Virtual Operator (MNVO) provides services to its customers who do not own any wireless infrastructure. A wireless virtualization enables MNVO to use multiple virtual wireless networks to provide services over the same physical infrastructure. The allocation of resources owned by the MNVO to its mobile customers is a challenging problem. The objective of MNVOs is to allocate resources (sub-channel and power) to maximize their respective revenues. Classical auction methods typically maximize social welfare and may often perform poorly in terms of revenues. In [39], the authors present a deep learning network to learn an auction-based method to allocate resources to the users of the MNVo. Fig. 21 shows the proposed auction model with the MNVO auctioning the resources (sub-channel and power) to a set of $N=\{1,\dots,n\}$ users. Let user $i$ get power $p_{i}$ and sub-channel $sc_{i}$. If these values are higher than the corresponding values for other users $j\neq i$, the valuation of the $i^{th}$ user is given as $v_{i}=p_{i}\times sc_{i}$. Let the payment for user $i$ be $t_{i}$, then the utility is
Figure 21: Auction for allocating sub-channels and power (Figure reproduced from [39])
@@
The optimization problem of Eqs. 98-100 is solved using MyersonNet [14] given in Figs. 5,6,7. It is observed that the deep learning based resource allocation improves the revenue for MNVO as compared to Second Price Auction (with traditional optimization).
@@ VIII-C A Procurement Auction for Agricultural Inputs
[4] extends the RegretNet architecture [14] to tackle the problem of volume discount procurement auctions in the agricultural setting. Farmer Producer Organisations (FPOs) or Farmer Collectives (FCs) typically consist of around thousand farmers. Individual farmers may have requirements such as seeds, pesticides, fertilizers, etc., but may not be able to bargain the best deal from the suppliers if they approach the suppliers on their own. FPOs collect the requirements of all their farmers, organize all the requirements into clusters, and approach the suppliers with a bulk requirement. Due to the large volume of procurement, suppliers generally offer volume discounts, i.e., a higher percentage of discount on the original asking price as the number of units purchased increases. This is reflected in the bids submitted by the suppliers. In addition to the constraints satisfied by the RegretNet architecture, namely, IR, DSIC, and revenue maximization, the authors aim to satisfy certain additional constraints. Revenue maximization here becomes cost minimization, as it is a procurement auction. The additional constraints are fairness (envy-freeness), social welfare maximization, and business constraints (minimum and maximum number of winning suppliers, minimum amount of business to be given to each winning supplier, etc.). There is, however, some compromise to the social welfare maximization.
@@
The volume discount procurement auction setting of [4] is as follows. Consider the case of a single buyer (FPO) who wishes to purchase $m$ homogeneous units of a single item from $n$ different suppliers. Let $m$ be divided into $k$ brackets. For each bracket, the bid submitted by the supplier (which is the price per unit in that bracket) must be less than or equal to the bid submitted by the supplier for the previous bracket. Each supplier $i$ submits a bid vector $b^{(i)}=(b_{1}^{(i)},b_{2}^{(i)},...,b_{k}^{(i)})$. Given all the input bids $b=(b^{(1)},b^{(2)},...,b^{(n)})$, the neural network outputs the allocation vector $a(b)=(a_{1}(b),a_{2}(b),...,a_{n}(b))$ and the payment vector $p(b)=(p_{1}(b),p_{2}(b),...,p_{n}(b))$ where $a_{i}(b)$ is the number of units sold by supplier $i$ and $p_{i}(b)$ is the payment made to supplier $i$.
@@
Through experimental analysis, the authors of [4] show how a DL based procurement auction captures the essence of the agricultural market better than existing analytical methods. Even as it minimizes additional constraints such as envy, when compared to VCG, their model saves the FPO more money.
@@ IX Summary and Conclusions
Table II provides a summary of the more prominent DL architectures for auction design as well as their evaluation. The DL approaches, in current literature, used for designing auctions include MenuNet, Deep Interest Network, RochetNet, MoulineNet, RegretNet, PreferenceNet, Deep Neural Auction and the Reinforcement Learning setups. MenuNet [43] is based on the principle of taxation, which lets buyers make selections that yield an IC mechanism. In online auctions, the essential task of click through rate prediction is implemented using Deep Interest Network [49]. MoulineNet is used to determine agents’ preferred optimal facility location [17]. An extension of RegretNet is PreferenceNet [38] that includes buyers’ preferences in Auction design. The Reinforcement learning setup is based on a no-regret learning mechanism and in the Deep Neural Auction model [26], the positional effect of ordering of sellers based on some metric is removed. The summary of key approaches referenced in our paper are given in Table II. In this table, strategyproofness indicates the type of IC constraint the model aims to satisfy, efficiency indicates whether it is revenue maximizing or welfare maximizing, fairness indicates the model’s aim to minimize violations of IR, envy, etc. and other constraints refer to constraints such as budget balance that the model aims to satisfy.
@@
Declaration of Competing Interest
The authors confirm that there are no financial conflicts of interest or personal relationships that could influence the reported findings in this manuscript.
Ethics approval and consent to participate
Not applicable, as no ethical approvals or consent required for this article
Consent for publication
Not applicable, as no individual patient data or identifying information is included in this manuscript.
Availability of data and material
Not applicable, as no data is required for this study
Competing interests
The authors have disclosed no competing interests related to this work.
Funding
The first author V. Udaya sankar received funding from TARE (Teachers Associateship for Research Excellence) (File No: TAR/2021/000114) Grant of the Department of Science and Technology, Government of India.
Authors’ contributions
V. Udaya Sankar (VUS) initiated the study and drafted the initial manuscript. Vishisht, Mayank, Narahari all are responsible for providing crucial inputs and contributing to the major revision along with VUS.